OpenStudy (anonymous):

find the solutions to the equation below. Check all that apply. x^2+14x+40=0 A.X=-10 B.x=6 C.x=-8 D.x=-4 E.x=4 F. X=3

OpenStudy (anonymous):

\[(x+4) (x+10) = 0\]What do you think?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Not quite... From the factored out equation, the addition sign means the solutions will be negative...

OpenStudy (anonymous):

C or D are the only negative ones

OpenStudy (anonymous):

What about A?

OpenStudy (anonymous):

yup and a

OpenStudy (anonymous):

OKay, so what's the answer? Don't guess!

OpenStudy (anonymous):

let me factor it

OpenStudy (anonymous):

What two numbers when added together make 14 and when multiplied together make 40?

OpenStudy (anonymous):

Dude, it is already factored..... \[x^2+14x+40=0 \]\[(x+4)(x+10)=0 \]

OpenStudy (anonymous):

4 nd 10

OpenStudy (anonymous):

positive or negative? what letters?

OpenStudy (anonymous):

D nd A

OpenStudy (anonymous):

YES! CORRECT